OpenStudy (accidentalaichan):

2(x + 6) > 3x − 2 I just need someone to help me check my work.

OpenStudy (accidentalaichan):

\[2 (x + 6) \ge 3x - 2\] \[4 (x + 6) \ge 3x\] \[4(6) \ge 2 x\] \[24 \ge 2x\] \[12 \ge x\]

OpenStudy (anonymous):

Let me see your work so I can check if you did it right.